# Open Quantum System Approaches to Superconducting Qubits

### v1

@@ V.7.1 Input-output theory in CQED
$H\simeq H_{\rm s}+H_{\rm tml}+\int_{0}^{\infty}\operatorname{d}\omega\lambda(\omega_{r})(ab^{\dagger}_{\omega}+a^{\dagger}b_{\omega}),$ (94)
@@
$-i$ $\int_{0}^{\infty}\operatorname{d}\omega\sqrt{\frac{\omega}{4\pi cv}}e^{-i(\omega-\omega_{r})t}(b_{\rm R\omega}-b_{\rm L\omega})=-\omega_{r}Z_{\rm tml}\frac{C_{k}}{C_{r}}\sqrt{\frac{\omega_{r}C_{r}}{2}}a,$ (108)
@@
$V(x,t)\simeq\sqrt{\frac{\omega_{r}Z_{\rm tml}}{2}}\Big{[}e^{i\omega_{r}x/v-i\omega_{r}t}b_{\rm out}(t)+e^{-i\omega_{r}x/v-i\omega_{r}t}b_{\rm in}(t)+{\rm H.C.}\Big{]},$ (118)
@@ V.8 Stochastic Approach
V.8.1 Stochastic Schrödinger equation
@@
$i\operatorname{d}|\Psi(t)\rangle=\Big{(}H_{\rm s}+H_{\varepsilon}+\lambda W\Big{)}|\Psi(t)\rangle\operatorname{d}t,$ (119)
@@
V.8.2 Stochastic master equation
@@
$K_{\alpha}=\lambda^{2}\sum_{\beta}\int_{0}^{\infty}\operatorname{d}\tau C_{\alpha,\beta}(\tau)e^{-iH_{\rm s}\tau}V_{\beta}e^{iH_{\rm s}\tau},$ (144)
@@ V.8.3 Stochastic approach in superconducting qubits
$\operatorname{d}\rho=-\frac{k}{2}[\sigma_{z},[\sigma_{z},\rho]]\operatorname{d}t+\sqrt{k}(\sigma_{z}\rho+\rho\sigma_{z}-2\langle\sigma_{z}\rangle\rho)\operatorname{d}\mathcal{W},$ (154)
@@ VI Summary
In this study, we have discussed various open quantum system approaches to model superconducting qubit dynamics under environmental noise that are useful for qubit design, control, and readout. These methods are given in FIG. 8. Each presented method is suited for specific noise types and qubit-environment interactions, providing insights into qubit dynamics. We aimed to explain in detail how these methods specifically apply to superconducting qubits. These techniques can help to shed light on decoherence mechanisms and qubit dynamics in CQED and related architectures. They remain indispensable for quantum noise engineering and mitigation efforts for robust quantum information processing. Considering the wide range of theoretical and experimental studies in this field, we cannot claim to have considered all possible methods, but we believe that this work can provide a comprehensive overview of important methods that can be considered by researchers.

### v2

@@ V.7.1 Input-output theory in CQED
$H\simeq H_{\rm s}+H_{\rm tml}+\int_{0}^{\infty}\operatorname{d}\omega\lambda(\omega_{r})(ab^{\dagger}_{\omega}+a^{\dagger}b_{\omega}).$ (94)
@@
$-i$ $\int_{0}^{\infty}\operatorname{d}\omega\sqrt{\frac{\omega}{4\pi cv}}e^{-i(\omega-\omega_{r})t}(b_{\rm R\omega}-b_{\rm L\omega})=-\omega_{r}Z_{\rm tml}\frac{C_{k}}{C_{r}}\sqrt{\frac{\omega_{r}C_{r}}{2}}a.$ (108)
@@
$V(x,t)\simeq\sqrt{\frac{\omega_{r}Z_{\rm tml}}{2}}\Big{[}e^{i\omega_{r}x/v-i\omega_{r}t}b_{\rm out}(t)+e^{-i\omega_{r}x/v-i\omega_{r}t}b_{\rm in}(t)+{\rm H.C.}\Big{]}.$ (118)
@@
For instance, input-output theory is particularly relevant for modeling the readout of transmon or fluxonium qubits coupled to microwave resonators, as commonly implemented in CQED architectures.
@@
In the next section, we use a stochastic approach to describe the average behavior of a superconducting qubit in the presence of a random interaction with the environment.
@@ V.8 Stochastic Approach
V.8.1 Stochastic Schrödinger equation (SSE)
@@
$i\operatorname{d}|\Psi(t)\rangle=\Big{(}H_{\rm s}+H_{\varepsilon}+\lambda W\Big{)}|\Psi(t)\rangle\operatorname{d}t.$ (119)
@@
V.8.2 Stochastic master equation (SME)
@@
$K_{\alpha}=\lambda^{2}\sum_{\beta}\int_{0}^{\infty}\operatorname{d}\tau C_{\alpha,\beta}(\tau)e^{-iH_{\rm s}\tau}V_{\beta}e^{iH_{\rm s}\tau}.$ (144)
@@ V.8.3 Stochastic approach in superconducting qubits
$\operatorname{d}\rho=-\frac{k}{2}[\sigma_{z},[\sigma_{z},\rho]]\operatorname{d}t+\sqrt{k}(\sigma_{z}\rho+\rho\sigma_{z}-2\langle\sigma_{z}\rangle\rho)\operatorname{d}\mathcal{W}.$ (154)
@@ VI Summary
In this study, we have discussed various open quantum system approaches to model superconducting qubit dynamics under environmental noise that are useful for qubit design, control, and readout. These methods are given in FIG. 8. Each presented method is suited for specific noise types and qubit-environment interactions, providing insights into qubit dynamics. We aimed to explain in detail how these methods specifically apply to superconducting qubits. These techniques can help to shed light on decoherence mechanisms and qubit dynamics in CQED and related architectures. They remain indispensable for quantum noise engineering and mitigation efforts for robust quantum information processing. The choice of approach often depends on the specific characteristics of the qubit-environment interaction being studied: phenomenological Lindblad equations are useful for general descriptions of decay and dephasing, Bloch-Redfield connects these rates to microscopic noise spectra, non-Markovian methods are crucial for memory effects, Floquet theory for periodically driven systems, input-output theory for systems coupled to transmission lines and measurement apparatus, and stochastic methods for simulating individual quantum trajectories or specific noise models.
While this review does not detail specific experimental coherence enhancements resulting from each theoretical method, it underscores that a deep understanding of qubit-environment interactions, facilitated by the discussed open quantum system approaches, is a prerequisite for designing robust qubits and developing effective error mitigation and correction strategies that lead to such improvements. Developing different methods to investigate the dynamics of a superconducting qubit as an open quantum system is vital for informing these advancements. Considering the wide range of theoretical and experimental studies in this field, we cannot claim to have considered all possible methods, but we believe that this work can provide a comprehensive overview of important methods that can be considered by researchers.